Check the youn g book for emple 010 02 OOO 04110/10 Grade O Type here to search
Answer #1

Part A)-

As per given data-

The magnitude of force \propto speed2 and We know that F = ma than-

| F| = Kv2 where K = proportion constant and F = ma

ma = Kv2 and we know that a=dv/dt so,

m*(dv/dt) = Kv2

m(dv/dt) - Kv2 = 0 .... (1) differential euqation

Part B)-

if we solve this differential eqn further -

and we will cross-multiplication here-

m(dv /v2) = K*dt

If we take intigration both the sides than-

m (v-1/-1) = K*t

1/v = - (K/m)*t velocity in function of time.

Thankyou !!! Have Fun
Please Rate High.